OpenStudy (katherinesmith):

Write the equation of an ellipse with vertices (10, 0) and (-10, 0) and co-vertices (0, 2) and (0, -2).

OpenStudy (tkhunny):

It's an ellipse. It looks like this: \(\dfrac{(x-h)^{2}}{a^{2}}+\dfrac{(y-k)^{2}}{b^{2}} = 1\) Looking at all the vertices, it is apparent that the center is (0,0). \(\dfrac{x^{2}}{a^{2}}+\dfrac{y^{2}}{b^{2}} = 1\) Looking at (10,0) and (-10,0), then a = 10 \(\dfrac{x^{2}}{10^{2}}+\dfrac{y^{2}}{b^{2}} = 1\) Looking at (0,2) and (0,-2), then b = 2 \(\dfrac{x^{2}}{10^{2}}+\dfrac{y^{2}}{2^{2}} = 1\) One for free. You show us another one.

OpenStudy (katherinesmith):

goodness. okay hold on.

OpenStudy (katherinesmith):

well my next one is the equation of an ellipse centered at the origin with foci at (0, -3) and (0, 3) and a major axis of 10. which is different than that one.

OpenStudy (tkhunny):

Barely different. Go! Follow my model or invent your own.

OpenStudy (katherinesmith):

how is it barely different! :(

OpenStudy (tkhunny):

Follow my model. One thing at a time.

OpenStudy (katherinesmith):

how is the foci and major axis anything like your model.

OpenStudy (tkhunny):

Are you going to work on it or argue with me? What do you know and what do you need to know? Same thing.

OpenStudy (tkhunny):

Major axis is in the same direction as the foci. The major axis length is twice "a". \(c^{2} = a^{2} - b^{2}\) Plenty of information. Go!!